STAR Protected
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC minutes April 5, 2021
Present: Alexei, Hongwei, Gene, Turi and Flemming
Hardware:
- TPC running well.
- Used access to put pizo driver on remote PS
- Adjusted methane content.
Operation:
Sector 20 RDO 3 had to be masked out. Unfortunately, the crew did not reboot iTPC on run control, so there are ‘fake’ cluster in the data. The manual for masking out has been updated to remind people to get this done.
There is currently no code to perform masking in the offline code. i.e. skipping hits in masked sectors.
Calibrations:
Gene summarized impressions from S&C calibration discussion. There are small deviations in M**2 plot with newest calibration. The deviation seen corresponds to order 0.2% in momentum. Indicates imperfect calibrations in TPC and/or BTOF. What are requirements to physics analysis to be performed? PWG should define?
It is important to have clarifications if current alignment and calibrations is sufficient to do physics for at least some physics analysis, since redoing / improving alignment will take several months
Yuri iterated the deviation is ~40% (at what momentum?) of mass**2 width. Have requested plots from BTOF group to get additional finer divided information. Is not convinced its TPC; Needs additional info to separate. Example: plotting tray numbers rather than phi.
Babu has done first pass on run-20 9.2 GeV dynamic distortions. Run 19 is done.
Yuri is revisiting dE/dx for both 19 and 14.5 GeV. Not satisfied with aging effect treatment.
If I recall, there was an early fast drop after installation of iTPC reported by Yuri long time ago.
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC meeting May 12, 2021
Present: Alexei, Gene, Yuri, Hongwei and Flemming
Hardware:
Basically in good shape. Struggled with HV GUI’s likely due to ITD scanning during the access day. Sector 20 outer was repaired by Tonko and Mike, bad RORC card.
Has observed the part short of west filed cage.
Alexei has the calibrated NMR in house.
Operations:
The O+O running very clean. The 1.5mrad crossing angle and levelling at ~3-4kHz Mbias results in very little if any dynamic space charge.
This bodes well for the FF setting to be run in about a week.
For the central collisions to be run right after the RFF 400 M Mbias the Mbias rate will be possibly ~20kHz.
Calibrations:
Gene looked at h+/hi ratios vs. pt for the 19.6 Gev au+au and reports seeing sector to sector variations;
https://drupal.star.bnl.gov/STAR/blog/genevb/h-h-run-19-196-gev
These variations are at similar level as after run 12 alignment calibrations i.e little overall effect but sector to sector variations.
The sectors showing largest effect correlates with Zaochen's findings in mass2 splitting in BTOF
Thus the variation is most likely due to TPC calibrations.
Yuri commented that until we have the FF data its unlikely that underlying reason, and improved calibration can be made. It is also likely to be a large effort, not a few weeks.
Yuri reported that dE/dx calibration for 14.5 and 19.6 is completed
It was also reported to S&C management
https://drupal.star.bnl.gov/STAR/system/files/19%20and%2014p5%20GeV%20dEdx%20Calibration.pdf
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
- hardware status
Gene: humidity impacts seen on inner field cage currents. See image shown further down, below agenda. More on this topic in general in my blog.
- operations
- calibration status

Time | Talk | Presenter |
---|---|---|
14:00 | OO 200 GeV 2021 RF vs FF ( 00:20 ) 1 file |
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
- hardware update
- operations
- calibration updates
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC meeting April21 minutes
Present:
Alexei,Babu, Flemming,Gene,Hongwei,Yuri
TPC running well. The pizo driver for the laser failed , no power,was adjusted during quick ccess, and then rebooted today during access. Lost about one laser run during. Will investigate how remote reboot can be done. Such failure has happened 1-2 per run.
Flemming present a few more observation on the single bad pad hot due to values this seem to have pedestal subtracted incorrectly. This is intermittent and usually, but not always clears on next run. Also it only 5-10% of events with this failure. See https://drupal.star.bnl.gov/STAR/blog/videbaks/tpc-qa-2021
for more details
The 19.6 GeV calibration sample is complete. Includes dynamic space charge, and event by event t0 correction. EPD had time offsets calibrated/changed at some point in time during the 19.6 GeV data sets. This change is now in DB. The 14.5 GeV should be done by the weekend. Babu will start looking at the run 20 soon.
The sample(s) can now be checked for BTOF calibrations, and dE/dx.
If the mass square difference seen in east persists with the DbTpcMaker that was updated, there will be additional work to identify this issue. The TOF group will not sign off with a mass split of approx. 1%.
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
- hardware status
- operations
- calibration status
Time | Talk | Presenter |
---|---|---|
15:00 | Averaged STAR magnet current table ( 00:10 ) 0 files | |
15:00 | Averaged STAR magnet current table ( 00:10 ) 0 files | |
15:10 | 2nd round with Run XIX calibration production for 19GeV and 14p5GeV samples ( 00:20 ) 0 files | Yuri Fisyak |
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
- hardware status
- operations
- software
- big events
- update on test productions
Minutes:
hardware: There was a failure of LV for TPX. broken fan; replaced with spare available in IR.
Have ordered 3 more fans, special. 3 mo. delivery
After a massive trip, one HV section 20-4 itpc seem not to have been brought up and not noticed. Asked David
to check the readouts during this event.
We see very large events in run 19 reconstruction, and in current data taking
a) current activity is related to the beat-squeeze; discussed also at ops meeting. Under discussion what to happen
b) run-19 reco is related to some store with BMM running throughout some runs.
Discussion on TPC calibration. Good progress is being made. Additional discussions later this day.
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
-hardware
-operation
- test production progress
- AOB
Time | Talk | Presenter |
---|---|---|
14:00 | new schema for Inner Sector alignment ( 00:20 ) 1 file | Hongwei Ke |
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
TPC weekly
virtual , at 19:00 (GMT), duration : 00:00
- hardware status
- operations
- software
- cluster quantization
-update on test productions
TPC weekly
virtual , at 20:00 (GMT), duration : 00:00
No issue with TPC hardware
-Have adjusted methan content to make Vdrift more constatnt; there were several period in Feb with variations.
- powerdips and magnet trips during last weeks running.
The two runs in day 38 and 66 had outer sector 19 exhibit large excess of charge that extended for all pad and rows for about 150 timebins. See blog entries in https://drupal.star.bnl.gov/STAR/blog/videbaks/tpc-qa-2021 near top of page
Most likely a faulty GG driver board . Alexei suggested looking at signals from board during next maintanence day (unless it get really bad). Will talk to Tim on this.
Gene reported looking at laser un running 7.7 GeV for local clock vs. coll clock. Did not see any different between laser positions. Now looking at prompt hits for two conditions, (UPDATE: small shifts seen for the entire TPC by 14.1 ns, but not on the scale of 1 time bucket; documentation is here)
Yuri show plots
these are references for plots which we discussed today
1. Comparison RF and FF https://drupal.star.bnl.gov/STAR/system/files/TpcCSTransformations_0.pdf
-- the FF has only 34k events.Alignment same for two settings.
2. QA plots for RunXXI (express production) : https://www.star.bnl.gov/~fisyak/star/BES/
---
3. Lasers for different fields
-- no clear pattern. Sector 20 seems consistent for all fields. 20 has best lasers
4. Prompts for different VPD timing cuts
-- shows dependence on collisions timing (sum of VPD timing)
The last plots on meeting page only
FST purple cable length
Final summary
The new plot (first) has the proper range of cable:
Final proposal is to divide into two groups